OpenStudy (anonymous):

the 13th term of an arithmetic sequence is 69. The common difference is 7. What is the first term of the sequence? Show all work.

OpenStudy (anonymous):

What's the formula to find the nth-term?

OpenStudy (experimentx):

the nth tern is given by tn = a + (n-1)d your n=13, tn = 69, d = 7 or, 69 = a + 12*7 or a = 69 - 84 = -15